OpenStudy (anonymous):

What is the value of the expression when a = 2? how would i be able to answer this confused 5a^4 4 a exponent

OpenStudy (anonymous):

oh god! @tom982

OpenStudy (anonymous):

2*2*2*2=16*5=80

jimthompson5910 (jim_thompson5910):

5a^4 is really 5*(a^4) You don't need that parenthesis, but it helps separate what is really being raised to the power of 4.

jimthompson5910 (jim_thompson5910):

since a = 2, you replace all of 'a' with 2 to get 5*(a^4) = 5*(2^4)

jimthompson5910 (jim_thompson5910):

from there, you would use the order of operations (PEMDAS) to evaluate 5*(2^4) the E for exponent comes before M for multiplication in PEMDAS. So you do exponentiation first and then you multiply. The parenthesis helps reinforce what comes first.

jimthompson5910 (jim_thompson5910):

so you could easily have 5*2^4 and mean the same thing as 5*(2^4)

OpenStudy (anonymous):

oooh okay

OpenStudy (anonymous):

wow i actually got it this time lol